Table: Prices and Demand Quantity of Hats Price Demanded Der Hat $30

(Table: Prices and Demand) The New Orleans Saints have a monopoly on Saints logo hats. The marginal cost of producing a hat is $18. If the Saints increase the number of hats they sell from 4 to 5, marginal revenue is:

$8.

$20.

$22.

$12.
Answer #1

Answer : The answer is option D.

Total revenue = Price * Quantity

For 4 hats the total revenue = 22 * 4 = $88

For 5 hats the total revenue = 20 * 5 = $100

Marginal revenue = Total revenue of 5 hats - Total revenue of 4 hats = 100 - 88 = $12

Therefore, option D is correct.